### v2

# Who is driving the conversation? Analysing the nodality of British MPs and journalists on Twitter

Abstract
Who sets the policy agenda? In this paper, we explore the roles of policy actors in agenda setting by studying their relative influence in policy-related discussions. Our approach builds on “nodality” —a concept in political science that determines the capacity of an actor to share information and to be at the centre of information networks. We propose a novel methodology that quantifies the nodality of all individual actors in any conversation by analysing a comprehensive set of their centrality measures in the related information network. We combine this with the analysis of the activity time-series, of the related conversation (or topic), to demonstrate how nodality scores relate to the capacity to drive topic-related activity. Here we analyse policy-related discussions on X (previously Twitter) and quantify the nodality of two sets of actors in the UK political system —Members of Parliament (MPs) and accredited journalists - on four policy topics: The Russia-Ukraine War, the Cost-of-Living Crisis, Brexit and COVID-19. Our results show that the capacity to influence the activity related to a topic is significantly and positively associated with nodality. In particular, we identify two dimensions of nodality that drive the capacity to influence topic-related activity. The first is “active nodality”, which reflects the level of topic-related engagement an individual actor has on the platform. The second dimension is “inherent nodality” which is entirely independent of the platform and reflects the actor’s institutional position (such as an MP in a front-bench role, or a journalist’s position at a prominent media outlet).
1 Introduction
Much political science research has been dedicated to understanding the relative influences of policy actors on the public policy agenda. Some scholars focus on the media [McCombs and Shaw, 2005], some on policymakers [Berkowitz, 1992] and some on the complex dynamics between the two [Wilson and Kingdon, 1993, Baumgartner et al., 2006, Baumgartner et al., 2013]. This paper tackles the long-standing question of ‘who sets the political agenda?’ through the lens of nodality – the capacity to exchange information – which was identified by Hood as one of the four key ‘tools of government’ used to make policy [Hood, 1983] and further developed by Hood and Margetts [Hood and Margetts, 2008] and Margetts and John [Margetts and John, 2023]. Nodality used to be something that government – as a set of organisations – possessed almost by virtue of being government, as the “water mill” at the centre of society’s information networks [Margetts and John, 2023]. But in a digital era, where policy debates take place in more open and diffuse information environments, governmental actors face greater competition for nodality [Castelnovo and Sorrentino, 2021]. Widespread use of the internet means that non-governmental organisations or individuals such as journalists, politicians, public figures and even citizens themselves can acquire nodality, which can challenge the government’s capacity to capture public attention and drive policy and media agendas [Margetts et al., 2019, Howlett, 2009, Howlett, 2022]. Therefore, as policy-related conversations are increasingly taking place outside institutional boundaries, there is a constant jostling for position between both governmental and non-governmental actors [Shearing and Wood, 2003].
Framed within the context of policy discourse in the UK, we seek to analyse the nodality of different actors on online platforms to determine how they drive the competition to shape policy conversations. While our ultimate aim is to analyse the relative nodality of a more diverse set of actors, including governmental departments, NGOs, local councils and public figures, in this paper we first test our methodology on a subset of elite actors in the UK: Members of Parliament (MPs) and accredited journalists. We propose a new method to dynamically measure nodality on social media, where information spreads widely and interactions are instantaneous. Although our methodology here is applied to Twitter (now X), it can be extended to other platforms and across a wide range of actors. Similar works include the study of exposure to misinformation from political elites in the USA [Mosleh and Rand, 2022], or the spread of news of COVID-19 from sources with different reputations [Gravino et al., 2022].
Influence on social media has been researched extensively in the field of communications (see for example, [Dubois and Gaffney, 2014, Anger and Kittl, 2011, Bakshy et al., 2011, Mei et al., 2015, Riquelme and González-Cantergiani, 2016, Esteve Del Valle and Borge Bravo, 2018, Rezaie et al., 2020, Kolic et al., 2022, Flamino et al., 2023, Zhang et al., 2023]). On X (previously Twitter), as with other social media sites, we assume that an actor has an influence on a particular topic when their activity generates activity by others (tweeting, retweeting, liking, replying or quoting) on the same topic. Understanding actors’ influence on the policy and media agenda will in part relate to these dynamics in policy conversations, which can be studied using transfer entropy, thus allowing for temporal and predictive analysis [Schreiber, 2000, James et al., 2016, Shorten et al., 2021].
In this paper, we investigate the nodality of two types of actors – MPs and journalists – that both play key roles in information exchange and influence in a policy system. Political representatives have an obligation to feed the concerns of their constituents into policy debates, while journalists disseminate information from policy discussions out to the public and wider information environment. We are able to obtain data for a substantial proportion of both sets of actors, which includes 581 out of 650 MPs and 8606 out of 9575 journalists [^1] The missing users do not have a Twitter account or the handles given could not be found. The dataset was also limited by the Twitter API policy which limits the number of tweets that can be retrieved per user to around 3200.. This approach helps to address a methodological challenge faced by previous work on policy actors and social media, where the differential attention that actors receive relate to both institutional and individual-level characteristics. For institutional differences, MPs can hold ministerial or shadow ministerial positions, which means that they are more likely than backbench MPs to receive attention (through more followers, for example), while journalists working for mainstream news channels or newspapers are likely to receive more attention by virtue of their institutional homes, while independent journalists may struggle to receive attention. There are institutional constraints for both sets of actors on social media: ministers often need to follow a Government or party line, and journalists may need to follow guidelines of their publishers. For individual characteristics, some individual MPs may become active and influential on a certain topic, possibly by virtue of their constituency being differently affected by an issue. Individual journalists may acquire a social media presence over and above the reputation of the publications for which they work. Our approach allows us to develop measures of nodality across these ’institutional’ and ’individual’ dimensions, thereby understanding where influence relates to an actor’s institutional position, and where it relates to individual-level activity, reputation or characteristics. Developing this methodology may help in the analysis of social media activity more generally, such as understanding widely varying levels of hate speech and abuse directed at representatives [Vidgen et al., 2020].
Nodality has been studied theoretically [Howlett, 2009, Howlett, 2022, Castelnovo and Sorrentino, 2021] and quantitatively, using network science -centrality measures [Margetts and Escher, 2006, Petricek et al., 2006], but we believe that this is the first attempt to quantify nodality as a way to assess actors’ communication capacity across different institutions and policy conversations, and over time, going beyond mere information transmission. To do so, we decompose nodality into two key components: (i) inherent nodality, which is earned by virtue of one’s institutional role (such as a frontbench position) and (ii) active or “topic specific” nodality, which is earned through high levels of engagement on a particular issue. We do this using different centrality measures and a transfer entropy analysis to establish a connection between the two dimensions of nodality and an agent’s influence.
We test the hypothesis that that influence will vary across different types of actor and policy topics on social media platforms [Bright, 2016, Bright et al., 2020, Sprejer et al., 2022, Parmelee, 2013, Usherwood and Wright, 2017]. We do this by using data from social media conversations around four policy topics in the year 2022: (1) the Russia-Ukraine war (2) the challenging economic situation facing ordinary citizens in the UK (Cost-of-Living crisis) (3) Britain’s departure from the EU (Brexit) and (4) the aftermath of the COVID-19 pandemic. All four issues were contested in policy terms and related to an increasingly challenging economic situation, driven by Brexit [Diamond, 2023, Dhingra and Sampson, 2022], the pandemic [Blundell et al., 2022], the Russia-Ukraine war [Mbah and Wasum, 2022, Liadze et al., 2023] and the resulting energy crisis [Harari et al., 2022, Francis-Devine et al., 2022], as well as a “mini-budget” in the autumn which led to sharp and sustained rises in interest rates [Chadha, 2022, Tosun and Lucey, 2023, Li, 2023]. We analyse Twitter data obtained for 581 MPs and 8,606 journalists registered in the UK between January 14, 2022 and January 13, 2023. We classify tweets using a weak-supervision classifier [Ratner et al., 2019, Castro-Gonzalez et al., 2024] and construct a weighted directed graph for each topic, where actors are nodes, and any interaction is captured as an information edge. To obtain group-level influence measures, we employ Transfer Entropy analysis on their activity time series, while inherent and active nodalities as defined above are derived through a Principal Component Analysis (PCA). We hypothesise that we will be able to distinguish institutional nodality as transferable across various topics, while active nodality is specific to a particular topic.
In this way, we address our central question: Which sets of actors drive the policy conversation on a given topic – which are the most influential nodes? In order to answer this question, we additionally tackle a further methodological issue: can we separate the relative effects of each actor’s ’institutional’ and ’individual’ nodality, and how can we measure nodality on online conversational networks across these two dimensions? In Section 2, we present our primary findings, leading us to subsequent discussions and conclusions in Section 3. Methods are presented in Section 4. For more detailed information about our methods, please refer to Appendix 5.
2 Results
2.1 Granger causality analysis and qualitative analysis of conversation networks
Figure 1: First insights for the Cost of Living Crisis conversation. (a) Activity time series for both the journalists and the MPs groups. (b) The giant component of the interaction network between journalists and MPs. The size of a node corresponds to the weighted out-degree which reflects the number of times a user has been retweeted (replied to or mentioned) by others in the network. Nodes representing journalists are shown in grey, while MPs are represented using coloured nodes where the colour signifies the political party they are affiliated with (Conservative, Labour, other parties). Edges here take the colour of the target node. Prominent nodes with high out-degrees have been labelled to highlight actors who are integral to the discussion (as a result of being heavily retweeted, replied to and mentioned). (c) The activity time series of Zarah Sultana, MP.
We begin by applying a Granger causality test on the activity time-series of MPs and journalists (shown in Figure 1a for the Cost of Living Crisis topic). This is a preliminary step in obtaining a general understanding of whether MPs or journalists dominate and lead the conversation on any of the four topics. We consider multiple time lags in our analysis (1 to 7 days lag) and find that in the cases of COVID-19 and Brexit, activity from journalists consistently “Granger-cause” the activity of MPs, with a statistical significance of $p<0.01$ for all considered time-lags, whereas for the Cost of Living Crisis and the Russo-Ukrainian war discussions, result are less clear and not statistically significant [^2] Details of the analyses for all topics can be found in Appendix 5..
We now proceed with a more comprehensive approach where we analyse the conversation network qualitatively for each topic. To do this, we build a directed network based on interactions between actors through retweets, replies and mentions, i.e. links represent the passage of information from a source actor who is retweeted (replied to or mentioned) by a target actor. Figure 1b shows the conversation network for the Cost of Living Crisis topic. We rely on network-based metrics [^3] Centrality measures have been used extensively in the past to determine the influence of actors on online platforms [Dubois and Gaffney, 2014]. to determine how central the actors are to each conversation. In our analysis, we find that nodes with the highest outdegrees are MPs with important parliamentary positions [^4] Prominent nodes include Cabinet and Shadow Cabinet MPs during 2022-2023, such as Boris Johnson, Rishi Sunak, Liz Truss, Keir Starmer, Rachel Reeves and Ed Milliband, or political editors at influential UK newspapers, such as Pippa Crerar (from The Guardian), Steven Swindon (The Times), and Jim Pickard (Financial Times).. Rishi Sunak has the highest outdegree in the network which can be attributed to the two key roles he held during the year 2022-23 —the Chancellor of the Exchequer, which is a pivotal role in public spending, and eventually that of the Prime Minister of the UK. We find that Rachel Reeves, the Shadow Chancellor of the Exchequer, and Nadhim Zahawi, who was the Chancellor under Liz Truss’ premiership in 2022, are also among the prominent nodes in the Cost of Living Crisis network. At the same time, we observe an anomaly here. One of the other actors central to the conversation is Zarah Sultana, a Backbench Labour MP, who despite her less significant role in the parliament has a high outdegree in the cost of living crisis network —an effect of a strong engagement with journalists who interact with her frequently on the topic. We further note that her nodality on the cost of living crisis discussion is topic-specific, and her activity, as well as nodality, is limited on the other topics (Figure 1c). In contrast, actors who have senior positions within the Cabinet or shadow frontbench such as Boris Johnson, Rishi Sunak, Liz Truss and Keir Starmer consistently appear central in all four conversations. These findings align well with the two dimensions of nodality outlined in Section 1inherent or institutional nodality which can be attributed to one’s institutional role, and active nodality which can be acquired at an individual-level by being active on a specific topic (such as in the case of Zarah Sultana).
2.2 Quantifying the nodality of MPs and journalists
We now attempt to quantify the inherent and active nodalities of actors in each of the studied conversations. The details of the approach are outlined in Section 4.3, but principally, we construct two networks for each topic: the conversation network (i.e. topic-specific network) and the null network capturing user interactions outside of the topic of focus. We then obtain the centrality measures of actors on both networks to perform a principal component analysis (PCA) for each topic, to quantify the inherent and the active nodalities of the actors. Guided by past work [Petricek et al., 2006, Margetts and Escher, 2006], we use six standard network-metrics for the PCA. In addition, we introduce two new metrics: the funnel bandwidth and the amplification bandwidth of an actor. The funnel bandwidth of a node measures its interaction activity, weighted by its own follower count, i.e. a measure of the extent to which an actor facilitates the passage of information in the network. Amplification bandwidth, on the other hand, measures the exposure of information from a user in two hops. We measure this as the number of times a user $i$ interacts with others, weighted by the number of followers of each target user $j$. These new measures are based on the follower count of the actors which has been extensively used in the literature to examine the role of users in online discussions [Riquelme and González-Cantergiani, 2016]. We thus finalise eight network-metrics —shown in Table 4.5 —, which on their own and combined with each other capture different aspects of nodality. We now employ the metrics in a PCA to quantify the two dimensions of nodality.
Figure 2: Details of the PCA and the $k$-means clustering for the Cost of Living Crisis conversation. (a) Figure showing the PCA results. (b) The three clusters represented in the interaction network. Nodes are coloured based on the cluster they belong to in the PCA (representing their nodalities in comparison to the mean nodality of the population). The size of the nodes reflects their weighted outdegrees. (c) Schematic of the general PCA results for the four topics. Observe that it has a notable triangular structure. Regions of sparse data are characterised by low active nodality (and high or low inherent nodality). Actors with low inherent nodality and limited participation in the conversation have low centrality measures and thus are excluded as outliers while fitting the data into a normal distribution prior to the PCA transformation. This is also true for actors with high inherent nodality and low active nodality. However nodes with high inherent nodalities are excluded less because they are retweeted by others in the network (given their prominent institutional roles).
We begin by generating all possible combinations of three or more measures —which gives us 255 combinations —of which we retain only those combinations that capture the inherent and the active nodalities of an actor along the first and second principal components. We verify this by employing an eigenvector test [^5] Details can be found in 4.3., where we ensure that the features from both networks either all positively (or negatively) correlate with the first principal component [^6] The weights of all features on the first principal component are either all positive or all negative.. Given that the inherent nodality of an actor is a reflection of their institutional role in the real-world and is independent of their activity on any topic, we hypothesise that the correlations in node measures on both the topic and the null networks as observed along the first principal component should capture the inherent nodality of an actor. We further assume that the residual nodality captured along the second principal component is the topic-specific active nodality. We verify this by ensuring that the features related to the topic network are positively correlated with the second principal component and orthogonal to the features obtained on the null network. In this way, we use the two principal components to distinguish between the inherent and active nodalities of actors. We only retain combinations of measures that satisfy the above form and we move into the next phase of our analysis where we identify the set of measures that maximises the number of common inherently nodal actors across all four topics. We expect actors with high inherent nodalities to consistently appear nodal, irrespective of the topic of focus. Therefore for each of the remaining set of features, we apply a $k$-means++ clustering, where $k=3$. The clustering approach groups actors based on their nodalities and we obtain three classes of policy-actors, those with inherent nodality: (i) higher than the mean of the studied population (nodal actors), (ii) same as the mean of the population and (iii) lower than the mean of the population (non-nodal actors). We then compare the nodal clusters across all four topics. The combination of metrics that maximises the number of shared inherently nodal actors across all topics includes: the strength of a node, their degree centrality, and the funnel bandwidth of the node.
Figure 3: Results of the PCA and $k$-means clustering for the four conversations. (a) The PCA results for the four topics. In a clockwise sense, starting from the top left, we find the results for the Russia-Ukraine War, Covid, Brexit and Cost of Living Crisis. (b) The result of the Transfer Entropy Analysis done for the resulting three clusters of the four conversations.
We now combine the clustering results from the PCA with the conversation network for the cost of living crisis discussion in Figure 2b. We find that the clusters form three tiers of the conversation network. Nodes in the inherently nodal cluster [^7] These are nodes with nodality score higher than the mean of the studied population. assume central positions in the conversation, i.e. we find that most of the interactions in the network are centred around these nodes and they exhibit high out-degree to in-degree ratios. We term these nodes leaders. Nodes with less than average inherent nodality (or the non-nodal actors) largely appear on the periphery of the network with limited connections to the rest of the network (average degree of 2). As these nodes typically appear to be at the periphery or the receiving end of the information propagation process, we term them receivers. Finally, we observe that nodes with average nodality form the layer between the central nodal cluster and the peripheral non-nodal cluster. We find that these nodes appear to bridge the interaction between the leaders and the receivers, by receiving information from the leaders (through retweets/mentions and replies) and passing it onto the receivers, and hence we term them funnelers.
To validate the PCA results and our hypothesis related to the role of each cluster in the topic-related discussion, we use a transfer entropy (TE) analysis. The TE analysis helps us measure the extent to which a group’s activity can predict the activity of others, measured as the reduction in the entropy of a group’s activity that can be estimated from the activity time series of the complement group [^8] Details of the method can be found in Section 4.4.. The difference in the share of entropy of Y predicted by the time series of X and the share of entropy of X predicted by the time series of Y, determines the direction of the flow of information. If this quantity is positive, there is an information flow from $X\rightarrow Y$, whereas if this is negative, information flows from $Y\rightarrow X$.
The results of the TE analysis aggregated across all four topics are shown in Figure 3. We find that leaders exhibit the highest share of influence among the three clusters. This supports our hypothesis that information in the network is typically disseminated from these nodes. Receivers, on the other hand, maintain a negative share of influence, which indicates that they mainly respond to the ongoing activity in the network. Whereas funnelers demonstrate low positive values suggesting that while their activity may influence the activity of the receivers, their own activity is influenced by the activity of the leaders. This means that information in the network flows from leaders to funnelers to receivers, as also concluded from our study of the topology of the network and the PCA. It is worth noting that despite of using two different data sets —conversation networks for the PCA and time series data for the TE analysis —our results from both analyses complement each other well.
2.3 Quantifying nodality at group level
Figure 4: Boxplots illustrating the inherent and active nodality scores of the different groups of MPs and journalists across all four topics, as obtained from the PCA. (a) shows the inherent nodality scores of each group while (b) shows the active nodality scores.
We now use the approach proposed above to answer our original research question —who among the MPs and journalists leads the conversation on the selected policy-related topics? So far we have shown that not all MPs (or journalists) have the same nodality, and that an actor’s capacity to influence online conversations often relies on their institutional roles. We therefore split MPs and journalists into sub-groups that reflect their organisational roles in the real world [^9] For the MPs, we exploit the structure of the British parliament to classify them into four groups: the Cabinet, which forms the executive power; the Shadow Cabinet, formed by the Opposition party that oversees the activities of the Cabinet; the Backbench, formed by MPs of the ruling party who do not have a place in the Cabinet; and its analogous for the opposition party, simply called the Opposition. We also divide journalists into two groups based on their follower counts. We find that the ratio of the Cabinet to the Backbench MPs and of the Shadow Cabinet to the Opposition (backbench) MPs is roughly 10%. Therefore we split the journalists by separating the 10% most-followed journalists from the rest of the journalists (and term them prominent journalists).. The groups are listed in Table 1, along with their sizes.
As before, we perform a PCA on the conversation networks for all four topics and we determine the mean inherent and active nodality score of each group. Results of the PCA are shown in Figure 4. We find that the Cabinet and the Shadow Cabinet MPs exhibit the highest inherent nodalities across all topics, typically followed by the prominent journalists, the backbench MPs (both government and opposition) and finally, by the rest of the journalists who exhibit the lowest inherent nodality scores. On examining the active nodalities of each group, we find that overall the median active nodality of all six groups of actors is similar to the mean active nodality of the studied population across all topics. Only the Shadow Cabinet distinguishes itself from the rest, by exhibiting an above average active nodality for the topic of Cost of Living crisis and a less than average active nodality for the Brexit-related conversation.
We simultaneously perform the TE analysis for each group and each topic over time. Results are shown in Table 1. For each topic, we highlight the top two groups with the highest shares of influence on the rest of the population [^10] A schematic of how this table is computed is shown in Figure 6 of Section 4.4.. We find that while the Cabinet MPs and the prominent journalists lead the activity on the Russo-Ukrainian conflict over time, the Shadow Cabinet and the Opposition backbench drive the discussion on the cost of living crisis. Furthermore, we observe that the Opposition backbench also leads the conversation on Brexit and COVID-19, along with prominent journalists.
Table 1: Temporal average influence score of each one of the groups for the different studied topics. Highlighted cells are those having a positive influence within the 95% confidence interval
Group Russia-Ukraine War COVID Cost of Living Crisis Brexit Number of nodes
Cabinet MP $6.53\%\pm 2.81\%$ $0.91\%\pm 2.45\%$ $0.02\%\pm 3.14\%$ $-1.41\%\pm 2.81\%$ 58
Backbench MP $1.11\%\pm 1.37\%$ $0.33\%\pm 1.10\%$ $0.10\%\pm 1.26\%$ $-1.37\%\pm 1.51\%$ 237
Shadow Cabinet MP $-0.1\%\pm 4.84\%$ $0.71\%\pm 4.25\%$ $9.41\%\pm 5.04\%$ $-1.08\%\pm 5.78\%$ 27
Opposition backbench MP $1.26\%\pm 0.99\%$ $1.11\%\pm 0.76\%$ $3.22\%\pm 1.41\%$ $1.85\%\pm 0.90\%$ 246
Prominent journalists (10% most followed) $2.75\%\pm 1.01\%$ $0.95\%\pm 0.79\%$ $0.38\%\pm 1.20\%$ $1.44\%\pm 1.16\%$ 657
Rest of journalists $-0.34\%\pm 0.94\%$ $-0.17\%\pm 0.77\%$ $-1.18\%\pm 1.48\%$ $-1.02\%\pm 1.63\%$ 3724
These results suggest that the ability of an actor to influence the discussion on a topic does not entirely depend on their institutional status but can also depend on other factors (such as their activity on the topic). To examine this further, we perform a regression analysis that determines the relationship between the capacity to influence the conversation on a topic and the different dimensions of nodality (i.e. inherent or institutional and active or individual nodality). To do this, we employ a mixed-effects model where the dependent variable is the influence exerted by a group on the activity of a topic, and the independent variables include the collective inherent and active nodalities of the groups and a related crossed term (inherent-active). Given that the influence of a group on the remaining population is measured using the time-series of their activities, we also include time as a independent variable. A two-week time period is used to compute the influence of each group (the dependent variable) and the PCA-related terms (the independent variables). Details of the model are given in Section 4.5.
Table 2: Results of the Linear regression for the model relating the group influence and the Nodality of its integrants.
$N=1045,\ R^{2}=0.6137$
Variable Coefficient $p$-value
Inherent** $0.010\pm 0.001$ 0.000
Active** $0.008\pm 0.004$ 0.025
Inherent-Active interaction $-0.001\pm 0.001$ 0.111
Time** $-0.002\pm 0.000$ 0.000
Intercept** $0.088\pm 0.029$ 0.002
Results of the regression analysis are presented in Table 2. We make four key observations: (i) the capacity of a group to influence the discussion on a specific topic depends not only on their institutional status but also on how active they are on the topic; (ii) the capacity of a group to influence the topic-related activity is positively associated with the mean inherent and active nodalities of the group; (iii) we also verify the independence of the two dimensions of nodality (inherent and active), given that the crossed term (inherent-active) does not significantly contribute to overall capacity of the group to influence the activity related to a conversation; and finally, (iv) assuming that the mean nodality of a group is fixed, we observe that the influence exerted by the group on the activity of a topic decreases with time.
3 Discussion
To date, the concept of nodality has been studied largely from a theoretical perspective [Howlett, 2009, Castelnovo and Sorrentino, 2021, Howlett, 2022, Margetts and John, 2023], with very few attempts to quantify nodality in an empirical setting [Petricek et al., 2006, Margetts and Escher, 2006]. Here for the first time, we focus on computing nodality scores for individual actors in policy conversations. More specifically, for our two sets of elite actors (MPs and journalists), we distinguish and measure two dimensions of nodality; (i) topic-specific active nodality at the individual level, relating to the activity of an actor on a specific topic, and (ii) inherent nodality relating to the actor’s institutional role and status. Distinguishing between the two types of nodalities provides a useful way to determine the nodal advantage of actors in prominent roles, such as in the case of Ministers and political editors, who by virtue of their institutional status possess nodality which can be transferred across topics. Our empirical results show that while the inherent nodality of institutional groups (such as Cabinet Ministers and backbench MPs) is consistent across topics, the individual (active) nodality of actors vary with the topic of focus. The methodology described in the paper can be generalised to any set of topics. For the four topics studied here, we find that node strength, degree centrality and funnel bandwidth (which is a measure reflecting the extent to which a node can amplify existing information in the network) is the best combination of measures to decompose nodality into individual (active) and institutional (inherent) nodalities.
Our results are further consolidated using a linear model which validates the independence of the two dimensions of nodality and establishes a relationship between the collective (mean) nodality of a group of actors and their capacity to influence the conversation on any given topic. A transfer entropy analysis is used to measure the influence of a group, in relation to the rest of the population, on any topic. We find that the Cabinet MPs and the prominent journalists hold the highest shares of influence in the discussion relating to the Russia-Ukraine War. Whereas the discussion on the COVID-19 pandemic is led by the opposition backbench MPs and the prominent journalists (there is however notably less activity on this topic as compared to other topics through 2022). This pattern continues in the discussion relating to Brexit (also with limited activity during 2022), where the conversation is again dominated by prominent journalists and opposition backbench MPs, with restrained activity on the topic from Labour and Conservative MPs in prominent roles. Finally, for the Cost-of-Living Crisis, we observe that the opposition MPs (both the shadow cabinet as well as the back-benchers) lead the conversation, highlighting reluctance of the ruling party to engage on this topic.
The main contribution of this work is to propose a method that measures the nodality of individual actors and their collective capacity to drive a conversation. For any given topic, our method measures the capacity of a group to influence the rest of the studied population, at any given point in time. In addition, by determining the scores for both individual and institutional nodality for each member of the group it also offers insights into the dynamics of influence within the group —enabling an understanding of the respective contributions of members (and of their role and activity) within the larger context. This work also enriches our understanding of the role of policy-actors in policy-related conversations, as it goes beyond studies which look at the influence of individuals in a conversation only in terms of their network attributes or positions in the network (i.e. centrality metrics) [Esteve Del Valle and Borge Bravo, 2018, Dubois and Gaffney, 2014, Riquelme and González-Cantergiani, 2016, Zhang et al., 2023]. Our study expands on the literature of politics and journalists on Twitter [Parmelee, 2013, Jungherr, 2014, Mosleh and Rand, 2022, Kim and Park, 2012, Graham et al., 2013, Graham et al., 2015, Bovet and Makse, 2019, Nożewski and Baranowski, 2021, Shapiro and Hemphill, 2017, Bright, 2018], introducing the actor’s role in their institutions and their individual activity on a topic as key elements in understanding the influence they exert.
Going forward, our methodology could support further research investigating the activities and influence of politicians on online platforms. For instance, here we classify inherent nodality as pertaining to institutional status, but there are other potential drivers of inherent nodality that our approach might help to identify, such as educational background, based on the hypothesis that elite actors attending the same private school or university are likely to be more connected in an online social network (as compared to a null network) —which might then drive their nodality on the platform, independent of individual levels of activity. Alternatively, our methodology might assist the study of different, more negative kinds of attention, such as hate speech directed at political representatives, where our approach might help assess which element of the hate directed at ministers relates to their institutional role, and which to their individual characteristics. This has been a challenge to examine in the past [Gorrell et al., 2020], given that ministers are more well-known and hence get more hate, regardless of race or gender [Agarwal et al., 2021]. Likewise, for journalists, who are increasingly suffering threats of violence online, such an approach might help determine whether the hate they receive relates to the publications they work for, or to their historic activity on policy-topics such as gender.
Finally, there are some limitations to this work. First, our elite networks include only British journalists and MPs. This is a limited set of individual actors, albeit an important one in terms of nodes that have both individual and institutional nodality, and who play an important role in setting the policy agenda. Second, we have covered only four policy topics, and future work should expand the range to further validate the analysis. In this way we might identify larger domains, such as defence and security (rather than the Ukraine war as focused on here), which could allow us to say something generalisable about elite networks in particular domains. Third, our results represent an analysis of the volume of activity – not the content. To tackle some of the methodological extensions noted above, it could be important to focus on content, such as different types of hate speech by targeted group, which have been shown to offer distinct challenges for classification and detection. Fourth, we have looked at only one specific social media platform - Twitter, now X, a platform currently experiencing a decline [^11] https://variety.com/2023/digital/news/musk-twitter-x-acquisition-one-year-user-revenue-decline-1235770297/, which may not be possible to analyse in this way in the future, due to the lack of data and also lack of universality (over the last years, virtually all politicians and journalists have maintained official Twitter accounts, but with the increasing uncertainty over X, they have started to transfer their activity to Instagram or Threads). It will be increasingly important therefore, to conduct such research on different platforms. Each one of the four limitations above could be tackled in future work: the expansion of types of actors, topics and online platforms.
4 Methods
4.1 Data set
Our database comprises of tweets published by British MPs and journalists between January 14, 2022 and January 14, 2023, collected using the Twitter API. Of the 650 House of Commons MPs, we find that 581 MPs have active Twitter accounts and a total of 580,312 tweets were collected from these users for the above time frame. We obtain a list of 9594 journalists from journalism.co.uk which lists journalists in the UK along with their Twitter accounts. Of these, we were able to retain information on 8606 journalists, and a total of about 5M tweets were collected over the given time period.
4.2 Dimensions of nodality
In [Petricek et al., 2006, Margetts and Escher, 2006], the authors present a qualitative analysis of multiple dimensions of nodality (for instance, extroversion, visibility of government websites in different countries) using network metrics such as centrality measures. Here we extend this analysis by proposing a methodology to quantify nodality on information networks, in addition to extending the dimensions of nodality studied in [Margetts and John, 2023]. We consider not only the extroversion and visibility of nodes in the information network, but also the within-node importance, the engagement a node generates, its authority, its capacity to transmit information, and its ability to amplify messages in the network. These additional dimensions are each paired with centrality measures, such as degree centrality, eigenvalue centrality, betweenness centrality, strength, and the authority and hubs measurements [Kleinberg, 1999]. These measures have been extensively used in the literature to gauge node influence in networks [Anger and Kittl, 2011, Bakshy et al., 2011, Mei et al., 2015, Riquelme and González-Cantergiani, 2016, Essaidi et al., 2020, Rezaie et al., 2020, Zhang et al., 2023]. A description of each dimension of nodality and how it applies to our research can be read in Table 4.5.
We also introduce two new measures which focus on the broad audience reached by individual actors in the network. These measures incorporate the number of followers of each node, also referred to as a measure of popularity of a node [Riquelme and González-Cantergiani, 2016], allowing us to capture the reach of a node in the network. These are:
$\nu^{i}=\sum_{h}\frac{f^{i}w_{hi}}{\langle k_{\mathrm{in}}\rangle^{*}}\hskip 10.0pt\mathrm{and}\hskip 10.0pt\mu^{i}=\sum_{j}\frac{f^{j}w_{ij}}{\langle k_{\mathrm{out}}\rangle^{*}}.$ (1)
$f^{i}$ represents the number of followers of node $i$, $w_{ij}$ is the weight of link $i\to j$, thus the number of times node $j$ interacts with information from $i$, and $\langle k_{\mathrm{in/out}}\rangle^{*}$ is the weighted in/out-degree average of the network.
While $\nu^{i}$ is interpreted as the extensive exposure that a node $i$ gives to other nodes through interactions, $\mu^{i}$ represents the exposure that nodes interacting with node $i$ provide by virtue of their own follower count. We call these two measures funnel bandwidth and amplification bandwidth respectively. Unlike the initial six measures that solely integrate structural information from the network, these two additional metrics incorporate Twitter-specific properties into the network analysis.
4.3 Active and inherent nodality through Principal Component Analysis
We perform a PCA [Jolliffe and Cadima, 2016] with different tests and checks to ensure that the resulting first two components directly correspond to the inherent and the active nodalities. At a higher level, this series of steps to obtain individual scores related to the active and inherent nodalities of users can be described as follows:
  1. 1.
    The PCA is performed for each topic using the measures presented in Table 4.5 for the topic-specific network and the null network. By definition, users have inherent nodality across both networks (as it should be independent of topics), while their active nodality exists only on the topic-specific network.
  2. 2.
    However, given the measures presented in Table 4.5, we understand that we are only studying a subset of conversations. Thus, to find the best combination of measures for our subset of conversations, we perform different PCAs with different offsets for each one of the features. We analyse all different combinations for the 8 measures, thus performing $\sum_{k=0}^{7}\binom{8}{k}=255$ PCAs per topic network, where $k$ is the number of offsets and 8 refers to the total number of features we have.
  3. 3.
    We filter the combinations of features that allow us to capture the inherent and active nodalities along the principal components. To achieve this, we perform a test on the first two eigenvectors of each PCA transformation. We only retain those feature combinations where (1) the first eigenvector from the PCA transformation has the same sign for all features across both networks, and (2) the signs of the entries for the null network in the second eigenvector are opposite to those of the topic-specific network. As an illustrative example, if suppose, we perform a PCA with any 3 features of the ones listed in Table 4.5, we obtain the related eigenvectors from the PCA transformation as $\boldsymbol{e}=(f_{1}^{t},f_{2}^{t},f_{3}^{t},f_{1}^{0},f_{2}^{0},f_{3}^{0})$, where $f_{n}$ refers to any of the three features from Table 4.5 and the superscript $t/0$ refers to the network it belongs to (topic-specific/null). Representing the signs of the entries of the eigenvectors as 1 and -1, two eigenvectors passing our test would look like $\boldsymbol{e_{1}}=(1,1,1,1,1,1)$ and $\boldsymbol{e_{2}}=(-1,-1,-1,1,1,1)$. For this particular example, $\boldsymbol{e_{1}}$ captures correlated features across the topic and the null network along the first principal component, while $\boldsymbol{e_{2}}$ extracts the features that are related to the topic-specific network.
  4. 4.
    Following our hypothesis of two types of nodality, from the results of our PCA we can distinguish between three types of users: those having high inherent nodality, those having average inherent nodality, and having less than average inherent nodality. Following this assumption, we perform a $k$-means++ [Arthur and Vassilvitskii, 2007] algorithm with $k=3$, each cluster containing one of the three types of users just described. The robustness of the clusters is then proven by a number ($M=50$) of repetition to then form stable clusters. This enables us to eliminate the bias given by the initial position of the centroids [Kanungo et al., 2004, Nielsen and Nock, 2015, Bahmani et al., 2012].
  5. 5.
    Once having the three clusters for a given PCA, we are now in search of the best combination for the four different topics. From out of these three clusters, we could expect, following our definition, that the high-inherent cluster and the low-inherent nodal cluster share a high number of nodes throughout the four topics. Thus, as our last check to pick the best combination of features for our set of four conversations, we select the one that maximises the number of nodes shared in each respective cluster between the four topics.
4.4 Influence score and Transfer Entropy
Transfer entropy [Schreiber, 2000] is typically interpreted as the shared information between X’s past and Y’s present, while also taking into account Y’s past. In [James et al., 2016], the authors highlight instances where Transfer Entropy can overestimate information flow and underestimate influence between two timeseries. Keeping this in mind, and to avoid any such errors we first determine the transfer entropy from $X$ to $Y$, $T_{X\to Y}$ as a share of the entropy of $Y$ at time $t$ and time delay $k=1$ and then compute the difference between the transfer entropy from $X$ to $Y$ and the transfer entropy from $Y$ to $X$, to estimate the net influence score such as in:
$\varphi(X,Y,t,k=1)=\frac{T_{X\to Y}(t,k=1)}{H_{Y}(t)}-\frac{T_{Y\to X}(t,k=1)}{H_{X}(t)},$ (2)
where $H_{X}(t)$ is the entropy of $X$ at time $t$. The influence score $\varphi$ at time $t$ gives the net share of information between $X$ and $Y$ at time $t$. $\varphi$ lies on the interval [-1, 1], where -1 refers to $Y$ causing the total entropy of $X$ at time $t$ and 1 represents the opposite. The transfer entropy $T_{X\to Y}$ is computed in the same way as originally presented in [Schreiber, 2000].
In our study, we measure the influence score of actors using their activity time series. The “shared information” here is related to the frequency of user activity on Twitter around a particular topic and therefore, the TE analysis it helps us measure the influence that a user has on someone else’s activity on the given topic. More details on how the influence score can be read in Appendix 5.3.
4.5 Relation between influence score and active and inherent nodality
We are interested in determining the relationship between the capacity of actors to influence conversations and the inherent and active nodalities of these actors. We are inspired by [Margetts and John, 2023], who define “nodal effectors” as a type of nodal actors who assumes a central position in an information network and enable the dissemination of knowledge in the network, while also influencing or nudging the behaviours of listeners. Here we think of MPs and journalists as nodal effectors who can similarly influence the extent to which other agents in the network engage in a conversation related to a particular topic.
We use the methodology described in 4.4 to determine the influence scores of actors at the group level (such as the Cabinet and the Shadow Cabinet) and we relate it to the individual nodality scores obtained using the PCA in the following way:
$\varphi(G,\neg G,t)=\mathrm{constant}+\gamma t+\sum_{i\in G}\alpha_{\mathrm{inherent}}\mathrm{PC}1_{i}^{t}+\alpha_{\mathrm{active}}\mathrm{PC}2_{i}^{t}+\beta\mathrm{PC}1_{i}^{t}\times\mathrm{PC}2_{i}^{t},$ (3)
where $G$ refers to a given group, which in our case can be the Cabinet, the Shadow Cabinet, etc. $\neg G$ refers to the complement of $G$, thus all nodes not belonging to $G$. PC1 and PC2 refer to the first two components of the resulting PCA analysis outlined in Section 4.3. $\alpha_{\mathrm{inherent}}$, $\alpha_{\mathrm{active}}$ and $\beta$ are three independent coefficients. The term PC1$\times$×PC2 is included to capture any correlations between PC1 and PC2. We also include a temporal term $\gamma t$ to capture any correlation between the influence and time.
We fit Eq. (3) with the values of $\varphi$ for different groups and times. To obtain temporal values from the PCA, we create snapshots of networks constructed over shorter time-frames that use the same time period as the TE to compute the influence scores. Here TE computes the measure of influence the activity of a group (over a time-period of 2 weeks [t-1,t]) exerts on the activity of the rest of the population at (t+1) week. The networks for the PCA analysis are similarly generated using tweet data during the weeks (t-1) and t, to obtain the nodality scores.
The centrality measures used to study different aspects of nodality and their interpretation.
[…]
Eigenvalue centrality A high eigenvalue centrality is related to the probability that information will pass through a particular node. In this sense, a node with a respective high eigenvalue centrality is interpreted as a node whose activity will most probably be seen in the conversation network by the other actors. Visibility
Betweenness centrality Betweenness centrality is related to the likelihood that information will pass through a particular node. Thus, an actor who is important to deliver a message in the conversation network is interpreted to be one with a high betweenness centrality. Within-nodes importance
Strength The strength of a node takes into account the number of links with other nodes and the weights associated with them. Thus, a node with a small number of ties but with high weights can be as “strong” as one with a high number of ties with low weights. Engagement
Authority (HITS algorithm) A high authority score reflects, in our case, a node that has a high number of out-degree links, thus having a high number of other nodes interacting with it. The algorithm reflects in this sense the authority of the node in the network [Kleinberg, 1999]. Authority
Hubs (HITS algorithm) A high hubs score reflects, in contrast with the authority score, a node that has a high number of in-degree links, thus reflecting the node interacting with a high number of other nodes. The algorithm reflects in this sense how much the node acts as a hub of information within the network [Kleinberg, 1999]. Information transmission
Funnel bandwidth $\nu$ This term reflects how much amplification can a node give to other’s messages when interacting with them. By having an important number of followers, a node can give others visibility by interacting with them. The more interactions, the more visibility a node gives to others. The former is then acting as a funnel of information Information transmission
Amplification bandwidth $\mu$ The amplification bandwidth relates to who interacts with a given node. The more nodes interacting with you, the more frequent those interactions are, and the more followers your interlocutors have, the more amplified your activity will be. Amplification
Acknowledgments.
The authors would like to thank Angus R. Williams for providing the first list of Twitter handles associated with MPs. The authors would also like to thank Eirini Koutsouroupa for her support and effort to get this work published.
Funding Statement.
This work was supported by the Public Policy Programme of The Alan Turing Institute under the EPSRC grant EP/N510129/1. This work was supported by the Ecosystem Leadership Award under the EPSRC Grant EP/X03870X/1 & The Alan Turing Institute.
Data Availability Statement.
Twitter data is made available in accordance with Twitter’s terms of service. Tweet IDs and analysis code are available at https://github.com/Turing-Online-Safety-Codebase/nodality. The corresponding tweets can be downloaded using the official Twitter API (https://developer.twitter.com/en/docs/twitter-api)